# An Inexact Bregman Proximal Difference-of-Convex Algorithm with Two Types of Relative Stopping Criteria

### v1

@@ 4 Convergence analysis
$0\in\partial F(\widehat{\bm{x}})=\partial\left(\iota_{\mathcal{Q}}+P_{1}-P_{2}\right)(\widehat{\bm{x}})+\nabla f(\widehat{\bm{x}})\subseteq\partial\left(\iota_{\mathcal{Q}}+P_{1}\right)(\widehat{\bm{x}})-\partial P_{2}(\widehat{\bm{x}})+\nabla f(\widehat{\bm{x}}),$
where the first inclusion follows from the generalized Fermat’s rule [38, Theorem 10.1], the equality follows from [38, Exercise 8.8(c)], and the last inclusion follows from [27, Corollary 3.4]. Thus, in this paper, we say that $\bm{x}^{*}$ is a stationary point of problem (1.1) if $\bm{x}^{*}$ satisfies
$0\in\partial\left(\iota_{\mathcal{Q}}+P_{1}\right)(\bm{x}^{*})-\partial P_{2}(\bm{x}^{*})+\nabla f(\bm{x}^{*}).$
We denote the set of all stationary points of problem (1.1) by $\mathcal{S}$.
@@
This type of assumption is commonly adopted to establish the convergence of the Bregman-type method in the nonconvex setting (see, for example, [10, Assumption D] and [42, Assumption 4]). It can be satisfied by, for example, $\phi(\bm{x})=\frac{1}{2}\|\bm{x}\|^{2}$ and $\phi(\bm{x})=\frac{1}{2}\|\bm{x}\|^{2}+\frac{1}{2}\|A\bm{x}\|^{2}$, both of which will be used in our numerical section. With these additional assumptions, we are now able to establish the global subsequential convergence of our iBPDCA as follows.
@@
where the last inclusion can be verified by the definition of the $\varepsilon$-subdifferential of $\iota_{\mathcal{Q}}+P_{1}$ at ${\bm{x}}^{k_{i}+1}$. Note that the sequence $\{\bm{\xi}^{k_{i}}\}$ is bounded due to the continuity and convexity of $P_{2}$ and the boundedness of $\{{\bm{x}}^{k_{i}}\}$. Thus, by passing to a further subsequence if necessary, we may assume without loss of generality that $\bm{\xi}^{*}:=\lim\limits_{i\to\infty}\bm{\xi}^{k_{i}}$ exists and $\bm{\xi}^{*}\in\partial P_{2}(\bm{x}^{*})$ due to the closedness of $\partial P_{2}$. In addition, we have $\delta_{k}\to 0$ and $\|\Delta^{k}\|\to 0$ from Proposition 3.1(iv) for (SC1) and Proposition 3.2(iv) for (SC2). Then, passing to the limit in (4.3), and invoking (4.1), (4.2), $\|\Delta^{k}\|\to 0$, Assumption B with $\{\bm{x}^{k_{i}}\}$ and $\bm{x}^{*}$ contained in $\mathcal{N}_{f}$, and the outer semicontinuity of $\partial_{\delta_{k}}\left(\iota_{\mathcal{Q}}+P_{1}\right)$ (see, for example, [15, Proposition 4.1.1]) with $\delta_{k}\to 0$, we obtain that
@@ 4.1 Global convergence of the whole sequence
$=\lim_{k\to\infty}F({\bm{x}}^{k})+\sigma\gamma_{\max}\lim_{k\to\infty}\mathcal{D}_{\phi}({\bm{x}}^{k},\,\bm{x}^{k-1})=\lim_{k\to\infty}F({\bm{x}}^{k})=\zeta.$
Statement (ii). We first note from the boundedness of $\{{\bm{x}}^{k}\}$ that $\emptyset\neq\Omega\subseteq\mathcal{S}$, where $\mathcal{S}$ is the set of all stationary points of problem (1.1). Take any $\bm{x}^{\infty}\in\Omega$ and let $\{{\bm{x}}^{k_{i}}\}$ be a convergent subsequence such that $\lim\limits_{i\to\infty}{\bm{x}}^{k_{i}}=\bm{x}^{\infty}$. Since $\bm{x}^{k_{i}}\in\mathrm{dom}\,P_{1}\cap\mathrm{int}\,\mathrm{dom}\,\phi$ for all $k_{i}\geq 0$, $P_{1}$ is a proper closed convex function and $\overline{\mathrm{dom}\,\phi}=Q$ (by Assumption A1), then $\bm{x}^{\infty}\in\mathrm{dom}\,P_{1}\cap\mathcal{Q}$, and $\nabla f(\bm{x}^{\infty})$ and $\nabla\phi(\bm{x}^{\infty})$ are well-defined from Assumptions B and C2, respectively. Moreover, it is easy to see that $\bm{x}^{k_{i}},\,\bm{x}^{\infty}\in\mathcal{X}\cap\mathcal{N}_{\phi}\subseteq\mathrm{int}\,\mathrm{dom}\,\phi$, where the open sets $\mathcal{X}$ and $\mathcal{N}_{\phi}$ are given in Assumptions A and C.
@@
$\langle\nabla f(\bm{x}^{k_{i}-1})+\gamma_{k_{i}-1}\big{(}\nabla\phi(\bm{x}^{k_{i}})-\nabla\phi(\bm{x}^{k_{i}-1})\big{)}-\Delta^{k_{i}-1},\,{\bm{x}}^{k_{i}}-\bm{x}^{\infty}\rangle\to 0.$
In addition, note that the sequence $\{\bm{\xi}^{k_{i}-1}\}$ is bounded and $P_{2}(\bm{x}^{\infty})-P_{2}({\bm{x}}^{k_{i}-1})\to 0$ due to the continuity and convexity of $P_{2}$. Using these facts and invoking $\delta_{k}\to 0$, we have that
@@
Assume that there exists a $K_{0}>0$ such that $\delta_{k}\equiv 0$ for all $k\geq K_{0}$.
Assumptions D1 and D2 are common technical assumptions employed in the global convergence analysis of existing Bregman-type algorithms or proximal DC-type algorithms; see, for example, [10, 42, 47]. Assumption E enforces that no error is allowed in the computation of $\partial P_{1}$ after finitely many iterations. This requirement is indeed achievable in many application problems; see examples in [44, 56] and the $\ell_{1-2}$ regularized least squares problem in our numerical section 5.1. But, it may also limit the theoretical understanding of our iBPDCA when it tackles a problem with complex constraints, for example, the constrained $\ell_{1-2}$ sparse optimization problem in our numerical section 5.2. It would be interesting to explore the possibility of eliminating such a requirement when establishing the global sequential convergence. We will leave it as a future research topic.
@@
Suppose that Assumptions A, B, C, D, E hold, $L<\gamma_{\min}\leq\gamma_{k}\leq\gamma_{\max}<\infty$, $0\leq\sigma<(\gamma_{\min}-L)/\gamma_{\min}$ for all $k\geq 0$, and $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$. Let $\{\bm{x}^{k}\}$ be the sequence generated by the iBPDCA with (SC1) in Algorithm 1. If $F$ is a KL function, then the sequence $\{\bm{x}^{k}\}$ converges to a stationary point of problem (1.1).
@@
In view of Theorem 4.1, we have that $\lim\limits_{k\to\infty}\mathrm{dist}({\bm{x}}^{k},\,\Omega)=0$ with $\Omega\subseteq\mathcal{S}$, where $\Omega$ is the set of all cluster points of $\{\bm{x}^{k}\}$ and $\mathcal{S}$ is the set of all stationary points of problem (1.1). Thus, we only need to show that the sequence is convergent.
@@
where the equality follows from $\bm{x}^{k}\in\mathrm{int}\,\mathrm{dom}\,\phi=\mathrm{int}\,\mathcal{Q}$ and hence $\partial\iota_{\mathcal{Q}}({\bm{x}}^{k})=\{0\}$ for all $k\geq 0$, and the last inclusion can be easily verified by the definition of the subdifferential of a proper closed convex function (see, for example, [37, Theorem 23.8]). Using the above two relations, one can obtain that, for any $k\geq\max\{K_{0},K_{1}\}+1$,
@@
We next show that the sequence $\{\bm{x}^{k}\}$ generated by the iBPDCA with (SC2) is also convergent to a stationary point of problem (1.1) under additional appropriate assumptions. The analysis follows a similar line of arguments as presented in Theorem 4.2 for the iBPDCA with (SC1), but will utilize the following auxiliary function with $\tau>0$:
@@
Suppose that Assumptions A, B, C, D, E hold, $L<\gamma_{\min}\leq\gamma_{k}\leq\gamma_{\max}<\infty$, $0\leq\sigma<(\gamma_{\min}-L)/\gamma_{\min}$ for all $k\geq 0$, and $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$. Let $\{\bm{x}^{k}\}$ be the sequence generated by the iBPDCA with (SC2) in Algorithm 1. If $H_{\tau}$ is a KL function for any $\tau>0$, then the sequence $\{\bm{x}^{k}\}$ converges to a stationary point of problem (1.1) for sufficiently small $\sigma$.
@@
In view of Theorem 4.1, we have that $\lim\limits_{k\to\infty}\mathrm{dist}({\bm{x}}^{k},\,\Omega)=0$ with $\Omega\subseteq\mathcal{S}$, where $\Omega$ is the set of all cluster points of $\{\bm{x}^{k}\}$ and $\mathcal{S}$ is the set of all stationary points of problem (1.1). Thus, we only need to show that the sequence is convergent.
@@
This implies that, for sufficiently small $\sigma$, the squence $\left\{H_{\tau}({\bm{x}}^{k},{\bm{x}}^{k-1})\right\}$ is non-increasing after a finite number of iterations. Moreover, since $F$ is level-bounded (by Assumption A4), we have that $F^{*}:=\mathrm{\inf}\big{\{}F(\bm{x})\,|\,\bm{x}\in\mathcal{Q}\big{\}}>-\infty$ and hence $\big{\{}F(\bm{x}^{k})\big{\}}_{k=0}^{\infty}$ is bounded from below. This together with the nonnegativity of $\frac{\tau}{2}\|\bm{x}^{k}-\bm{x}^{k-1}\|^{2}$ implies that $\left\{H_{\tau}({\bm{x}}^{k},{\bm{x}}^{k-1})\right\}$ is also bounded from below. Then, we can conclude that $\lim\limits_{k\to\infty}H_{\tau}({\bm{x}}^{k},{\bm{x}}^{k-1})=\zeta$ exists. In the following, we consider two cases.
@@
Before closing this section, we would like to make some remarks regarding the global sequential convergence results. First, one may have noticed that an additional summable error condition $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$ is required in Theorems 4.2 and 4.3 to establish strong convergence results. This summable error condition can be trivially satisfied if, for instance, no errors occur on the left-hand side of condition (3.2) (namely, $\Delta^{k}\equiv 0$) when applying a certain subsolver for solving the subproblem (3.1). Moreover, it has been established that $\sum^{\infty}_{k=0}\mathcal{D}_{\phi}(\bm{x}^{k+1},\,\bm{x}^{k})<\infty$ in (3.9) and (3.13) for (SC1) and (SC2), respectively. Thus, one could also verify $\|\Delta^{k}\|\leq\sigma\gamma_{k}\mathcal{D}_{\phi}(\bm{x}^{k+1},\,\bm{x}^{k})$ along with (SC1) or verify $\|\Delta^{k}\|\leq\sigma\gamma_{k}\mathcal{D}_{\phi}(\bm{x}^{k},\,\bm{x}^{k-1})$ along with (SC2) to ensure $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$. Interestingly, we have observed from our numerical experiments that it is usually adequate to simply employ our stopping criterion (SC1) or (SC2) to obtain satisfactory empirical performances. Therefore, in our experiments, we choose not to verify additional inequality to ensure $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$ for the sake of simplicity. Of course, how to eliminate such a summable error condition in theory is also an interesting future research topic. Second, a similar auxiliary function in the form of $H_{\tau}$ has also been used in, for example, [47, 48], for developing the global sequential convergence under the KL property. When $F$ is semialgebraic, $H_{\tau}$ with any $\tau>0$ is also semialgebraic and hence is a KL function; see, for example, [3, Section 4.3] and [8, Section 2]. More discussions can be found in [47, Remark 4.2].

### v2

@@ 4 Convergence analysis
$0\in\partial F(\widehat{\bm{x}})=\partial\left(\iota_{\mathcal{Q}}+P_{1}-P_{2}\right)(\widehat{\bm{x}})+\nabla f(\widehat{\bm{x}})\subseteq\partial\left(\iota_{\mathcal{Q}}+P_{1}\right)(\widehat{\bm{x}})-\partial P_{2}(\widehat{\bm{x}})+\nabla f(\widehat{\bm{x}}),$ (4.1)
where the first inclusion follows from the generalized Fermat’s rule [41, Theorem 10.1], the equality follows from [41, Exercise 8.8(c)], and the last inclusion follows from [31, Corollary 3.4]. Thus, in this paper, we follow [50, Definition 4.1] to say that $\bm{x}^{*}$ is a stationary point of problem (1.1) if $\bm{x}^{*}$ satisfies
$0\in\partial\left(\iota_{\mathcal{Q}}+P_{1}\right)(\bm{x}^{*})-\partial P_{2}(\bm{x}^{*})+\nabla f(\bm{x}^{*}).$ (4.2)
Moreover, if $P_{2}$ is continuously differentiable near $\bm{x}^{*}$, the last inclusion in (4.1) can be shown to hold as an equality at $\bm{x}^{*}$. In this case, we follow [13, Section 3] to say that $\bm{x}^{*}$ is an $\ell$(imiting) stationary point of problem (1.1) if $\bm{x}^{*}$ satisfies
$0\in\partial\left(\iota_{\mathcal{Q}}+P_{1}\right)(\bm{x}^{*})-\nabla P_{2}(\bm{x}^{*})+\nabla f(\bm{x}^{*}).$ (4.3)
Note that, an $\ell$-stationary point, as defined in (4.3), can also be shown to be a $d$(irectional)-stationary point; see, for example, [29, Proposition 1] and [30, Proposition 1]. In the following, we denote the set of all stationary points of problem (1.1) by $\mathcal{S}$.
@@
This type of assumption is commonly adopted to establish the convergence of the Bregman-type method in the nonconvex setting (see, for example, [10, Assumption D] and [45, Assumption 4]). It can be satisfied by, for example, $\phi(\bm{x})=\frac{1}{2}\|\bm{x}\|^{2}$ and $\phi(\bm{x})=\frac{1}{2}\|\bm{x}\|^{2}+\frac{1}{2}\|A\bm{x}\|^{2}$, both of which will be used in our numerical section. With these additional assumptions, we are now able to establish the global subsequential convergence of our iBPDCA as follows.
@@
where the last inclusion can be verified by the definition of the $\varepsilon$-subdifferential of $\iota_{\mathcal{Q}}+P_{1}$ at ${\bm{x}}^{k_{i}+1}$. Note that the sequence $\{\bm{\xi}^{k_{i}}\}$ is bounded due to the continuity and convexity of $P_{2}$ and the boundedness of $\{{\bm{x}}^{k_{i}}\}$. Thus, by passing to a further subsequence if necessary, we may assume without loss of generality that $\bm{\xi}^{*}:=\lim\limits_{i\to\infty}\bm{\xi}^{k_{i}}$ exists and $\bm{\xi}^{*}\in\partial P_{2}(\bm{x}^{*})$ due to the closedness of $\partial P_{2}$. In addition, we have $\delta_{k}\to 0$ and $\|\Delta^{k}\|\to 0$ from Proposition 3.1(iv) for (SC1) and Proposition 3.2(iv) for (SC2). Then, passing to the limit in (4.6), and invoking (4.4), (4.5), $\|\Delta^{k}\|\to 0$, Assumption B with $\{\bm{x}^{k_{i}}\}$ and $\bm{x}^{*}$ contained in $\mathcal{N}_{f}$, and the outer semicontinuity of $\partial_{\delta_{k}}\left(\iota_{\mathcal{Q}}+P_{1}\right)$ (see, for example, [18, Proposition 4.1.1]) with $\delta_{k}\to 0$, we obtain that
@@ 4.1 Global convergence of the whole sequence
$=\lim_{k\to\infty}F({\bm{x}}^{k})+\sigma\gamma_{\max}\lim_{k\to\infty}\mathcal{D}_{\phi}({\bm{x}}^{k},\,\bm{x}^{k-1})=\lim_{k\to\infty}F({\bm{x}}^{k}),$
which implies that $\lim\limits_{k\to\infty}F({\bm{x}}^{k})$ exists. For simplicity, we also use $\zeta$ to denote this limit in the rest of this paper.
Statement (ii). We first note from the boundedness of $\{{\bm{x}}^{k}\}$ that $\emptyset\neq\Omega\subseteq\mathcal{S}$, where $\mathcal{S}$ is the set of all stationary points of problem (1.1). Take any $\bm{x}^{\infty}\in\Omega$ and let $\{{\bm{x}}^{k_{i}}\}$ be a convergent subsequence such that $\lim\limits_{i\to\infty}{\bm{x}}^{k_{i}}=\bm{x}^{\infty}$. Since $\bm{x}^{k_{i}}\in\mathrm{dom}\,P_{1}\cap\mathrm{int}\,\mathrm{dom}\,\phi$ for all $k_{i}\geq 0$, $P_{1}$ is a proper closed convex function and $\overline{\mathrm{dom}\,\phi}=Q$ (by Assumption A1), then $\bm{x}^{\infty}\in\mathrm{dom}\,P_{1}\cap\mathcal{Q}$, and $\nabla f(\bm{x}^{\infty})$ and $\nabla\phi(\bm{x}^{\infty})$ are well-defined from Assumptions B and C2, respectively. Moreover, it is easy to see that $\bm{x}^{k_{i}},\,\bm{x}^{\infty}\in\mathcal{X}\cap\mathcal{N}_{\phi}\subseteq\mathrm{int}\,\mathrm{dom}\,\phi$, where the sets $\mathcal{X}$ and $\mathcal{N}_{\phi}$ are given in Assumptions A and C, respectively.
@@
$\quad-\langle\nabla f(\bm{x}^{k_{i}-1})-\nabla f(\bm{x}^{\infty})+\gamma_{k_{i}-1}\big{(}\nabla\phi(\bm{x}^{k_{i}})-\nabla\phi(\bm{x}^{k_{i}-1})\big{)}-\Delta^{k_{i}-1},\,{\bm{x}}^{k_{i}}-\bm{x}^{\infty}\rangle+\delta_{k_{i}-1}.$
@@
$\langle\nabla f(\bm{x}^{k_{i}-1})-\nabla f(\bm{x}^{\infty})+\gamma_{k_{i}-1}\big{(}\nabla\phi(\bm{x}^{k_{i}})-\nabla\phi(\bm{x}^{k_{i}-1})\big{)}-\Delta^{k_{i}-1},\,{\bm{x}}^{k_{i}}-\bm{x}^{\infty}\rangle\to 0.$
In addition, it follows from the continuity and convexity of $P_{2}$ and $\lim\limits_{i\to\infty}{\bm{x}}^{k_{i}-1}=\bm{x}^{\infty}$ (by (4.4)) that the sequence $\{\bm{\xi}^{k_{i}-1}\}$ is bounded, $\langle\bm{\xi}^{k_{i}-1},\,{\bm{x}}^{k_{i}-1}-{\bm{x}}^{\infty}\rangle\to 0$, and $P_{2}(\bm{x}^{\infty})-P_{2}({\bm{x}}^{k_{i}-1})\to 0$. Using these facts and invoking $\delta_{k}\to 0$, we have that
@@
Assumptions D1 and D2 are common technical assumptions employed in the global convergence analysis of existing Bregman-type algorithms or proximal DC-type algorithms; see, for example, [10, 45, 50].
@@
Assume that there exists a $K_{0}>0$ such that $\delta_{k}\equiv 0$ for all $k\geq K_{0}$ in (3.2).
Note that, under Assumption E, no error $\delta_{k}$ is allowed in the computation of $\partial P_{1}$ after finitely many iterations, while a non-zero error term $\Delta^{k}$ remains permissible in (3.2). Consequently, even with this assumption, Algorithm 1 retains its inexact nature compared to the exact BPDCA studied in [45], and thus the convergence of the whole sequence must still be established accordingly. While Assumption E looks restrictive, it is indeed achievable in many application problems; see examples in [47, 59] and the $\ell_{1-2}$ regularized least squares problem in our numerical section 5.1. Nevertheless, Assumption E could limit the applicability of our iBPDCA when addressing a problem with complex constraints, for example, the constrained $\ell_{1-2}$ sparse optimization problem in our numerical section 5.2. Investigating the possibility of eliminating this requirement when establishing the convergence of the whole sequence remains an interesting direction for our future research.
@@
Suppose that Assumptions A, B, C, D, E hold, $L<\gamma_{\min}\leq\gamma_{k}\leq\gamma_{\max}<\infty$, $0\leq\sigma<(\gamma_{\min}-L)/\gamma_{\min}$ for all $k\geq 0$, and $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$. Let $\{\bm{x}^{k}\}$ be the sequence generated by the iBPDCA with (SC1) in Algorithm 1. If $F$ is a KL function, then the sequence $\{\bm{x}^{k}\}$ converges to an $\ell$-stationary point of problem (1.1).
@@
In view of Theorem 4.1, we have that $\lim\limits_{k\to\infty}\mathrm{dist}({\bm{x}}^{k},\,\Omega)=0$ with $\Omega\subseteq\mathcal{S}$, where $\Omega$ is the set of all cluster points of $\{\bm{x}^{k}\}$ and $\mathcal{S}$ is the set of all stationary points of problem (1.1). Note that, under Assumption D2, a stationary point as defined in (4.2) is inherently an $\ell$-stationary point as defined in (4.3). Thus, we only need to show that the sequence is convergent.
@@
where the equality follows from $\bm{x}^{k}\in\mathrm{int}\,\mathrm{dom}\,\phi=\mathrm{int}\,\mathcal{Q}$ and hence $\partial\iota_{\mathcal{Q}}({\bm{x}}^{k})=\{0\}$ for all $k\geq 0$, and the last inclusion can be easily verified by the definition of the subdifferential of a proper closed convex function (see, for example, [40, Theorem 23.8]). Using the above two relations, one can obtain that, for any $k\geq\max\{K_{0},K_{1}\}+1$,
@@
We next show that the sequence $\{\bm{x}^{k}\}$ generated by the iBPDCA with (SC2) is also convergent to an $\ell$-stationary point of problem (1.1) under additional appropriate assumptions. The analysis follows a similar line of arguments as presented in Theorem 4.2 for the iBPDCA with (SC1), but will utilize the following auxiliary function with $\tau>0$:
@@
Suppose that Assumptions A, B, C, D, E hold, $L<\gamma_{\min}\leq\gamma_{k}\leq\gamma_{\max}<\infty$, $0\leq\sigma<(\gamma_{\min}-L)/\gamma_{\max}$ for all $k\geq 0$, and $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$. Let $\{\bm{x}^{k}\}$ be the sequence generated by the iBPDCA with (SC2) in Algorithm 1. If $H_{\tau}$ is a KL function for any $\tau>0$, then the sequence $\{\bm{x}^{k}\}$ converges to an $\ell$-stationary point of problem (1.1) for sufficiently small $\sigma$.
@@
In view of Theorem 4.1, we have that $\lim\limits_{k\to\infty}\mathrm{dist}({\bm{x}}^{k},\,\Omega)=0$ with $\Omega\subseteq\mathcal{S}$, where $\Omega$ is the set of all cluster points of $\{\bm{x}^{k}\}$ and $\mathcal{S}$ is the set of all stationary points of problem (1.1). Note that, under Assumption D2, a stationary point as defined in (4.2) is inherently an $\ell$-stationary point as defined in (4.3). Thus, we only need to show that the sequence is convergent.
@@
This implies that, for sufficiently small $\sigma$, the squence $\left\{H_{\tau}({\bm{x}}^{k},{\bm{x}}^{k-1})\right\}$ is non-increasing after a finite number of iterations. Moreover, since $F$ is level-bounded (by Assumption A4), we have that $F^{*}:=\mathrm{\inf}\big{\{}F(\bm{x})\,|\,\bm{x}\in\mathcal{Q}\big{\}}>-\infty$ and hence $\big{\{}F(\bm{x}^{k})\big{\}}_{k=0}^{\infty}$ is bounded from below. This together with the nonnegativity of $\frac{\tau}{2}\|\bm{x}^{k}-\bm{x}^{k-1}\|^{2}$ implies that $\left\{H_{\tau}({\bm{x}}^{k},{\bm{x}}^{k-1})\right\}$ is also bounded from below. And it follows from the strong convexity of $\phi$ on $\mathcal{N}_{\phi}$ (by Assumption C1) and Proposition 3.2(iii) that $\bm{x}^{k}-\bm{x}^{k-1}\to 0$. Then, combining with Proposition 4.1(i), we can conclude that $\lim\limits_{k\to\infty}H_{\tau}({\bm{x}}^{k},{\bm{x}}^{k-})=\lim\limits_{k\to\infty}F({\bm{x}}^{k})+\lim\limits_{k\to\infty}\frac{\tau}{2}\|{\bm{x}}^{k}-\bm{x}^{k-1}\|^{2}=\lim\limits_{k\to\infty}F({\bm{x}}^{k})=\zeta$. In the following, we consider two cases.
@@
Before closing this section, we would like to make some remarks regarding the global sequential convergence results. First, one may have noticed that our current analysis relies on the summable error condition $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$ to establish strong convergence results in Theorems 4.2 and 4.3. Fortunately, it has been shown that $\sum^{\infty}_{k=0}\mathcal{D}_{\phi}(\bm{x}^{k+1},\,\bm{x}^{k})<\infty$ in (3.9) and (3.13) for (SC1) and (SC2), respectively. Thus, one could also verify that $\|\Delta^{k}\|\leq\sigma\gamma_{k}\mathcal{D}_{\phi}(\bm{x}^{k+1},\,\bm{x}^{k})$ along with (SC1) or verify that $\|\Delta^{k}\|\leq\sigma\gamma_{k}\mathcal{D}_{\phi}(\bm{x}^{k},\,\bm{x}^{k-1})$ along with (SC2) to ensure $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$. Interestingly, our numerical experiments show that it is usually sufficient to solely employ our stopping criterion (SC1) or (SC2) for obtaining satisfactory empirical performance. Therefore, in our experiments, we choose not to verify the additional inequality to ensure $\sum_{k=0}^{\infty}\|\Delta^{k}\|<\infty$ for the sake of simplicity. Of course, eliminating the need for such a summable error condition in theory remains an interesting topic for future research. Second, a similar auxiliary function in the form of $H_{\tau}$ has also been used in, for example, [50, 51], for developing the global sequential convergence under the KL property. When $F$ is semialgebraic, $H_{\tau}$ with any $\tau>0$ is also semialgebraic and hence is a KL function; see, for example, [3, Section 4.3] and [8, Section 2]. More discussions can be found in [50, Remark 4.2].